Short Answer Question 2:

Company C has a $50M line of credit with a 5% compensating balance. The firm has outstanding borrowings of $37M. The fee on the unused portion of the line of credit is 0.25% and the interest rate on the borrowings is LIBOR of 0.28% plus 3.25%. What is the annual interest rate for the line of credit?
Answer #1

We are given,

Credit line = $50M

Outstanding amount = $37M

Compensating balance = 5%

Net disbursement = Outstanding amount - 5%*Credit line = 37 - 5%*5M = $34.5M

Interest rate(r) = 0.28% + 3.23% = 3.53%

Interest = r * Outstanding amount = 3.53% * 37 = 1.3061

Unused portion = 50 - 37 = $13M

Fee on unused portion = 0.25%

Fee amount = 0.25 * 13M = 0.0325M

Total interest and bank charges = 1.3061 + 0.0325 = 1.3386M

Effective annual interest rate = Total interest / Net disbursement = 1.3386 / 34.5 = 3.88%

Hence annual interest rate is 3.88%.

An upvote would be appreciated.